OpenStudy (jiteshmeghwal9):

A man buys eggs at 2 for Rs.1 and an equal number at 3 for Rs.2 and sells the whole at 5 for Rs.3.Find his gain or loss percent.

OpenStudy (jiteshmeghwal9):

@hartnn plz help :)

hartnn (hartnn):

'buys eggs at 2 for Rs.1' means ?

hartnn (hartnn):

at 2 means ?

hartnn (hartnn):

at 2 o'clock?

OpenStudy (jiteshmeghwal9):

nope but i also don't know :|

OpenStudy (jiteshmeghwal9):

i m thinking 'at 2 means' that he buys 2 eggs for Rs.1

OpenStudy (jiteshmeghwal9):

Daya kar ke javab de do :<

OpenStudy (jiteshmeghwal9):

@UnkleRhaukus plz help :)

ganeshie8 (ganeshie8):

X : 2eggs for Rs.1 => cost of each egg = Rs. 1/2 X : 3eggs for Rs.2 => cost of each egg = Rs. 2/3 total Cost of eggs = X[1/2] + X[2/3] = X [1/2 + 2/3 ] = X [7/6] = X[35/30]

ganeshie8 (ganeshie8):

Selling : 2X : 5eggs for Rs.3 => cost of each egg = Rs. 3/5 total Sales revenue = 2X[3/5] = X[6/5] = X[36/30]

ganeshie8 (ganeshie8):

Profit = Sales Revenue - Cost = X[36/30] - X[35/30] = X[1/30]

ganeshie8 (ganeshie8):

Profit % = (Profit/Cost)*100 = (X[1/30] / X[36/30]) * 100 = (1/36)*100 = ? %

OpenStudy (jiteshmeghwal9):

Thanx @ganeshie8 sir :)

ganeshie8 (ganeshie8):

yw friend :)

OpenStudy (jiteshmeghwal9):

:)